A constant retarding torque of 12Nm stops a rollin wheel of diameter 0.8m in a distance of 15m within 35s

a) Use work-energy theorem to calculate the initial angular speed when the torque was applied if the wheel's moment of inertia is 0.7kgm^2
Answer #1

angular displacement traveled

x = 15 / (2 pi 0.4) = 5.9683 rad

using work-energy theorem

T x = 0.5 I w^2

12* 5.9683 = 0.5* 0.7* w^2

w = 14.3 rad/s

====

Comment before rate in case any doubt, will reply for sure.. goodluck